OpenStudy (anonymous):

If a 30N net force acts on a 20kg object for 5s how fast will the object be moving at the time and how far will it have traveled?

OpenStudy (anonymous):

acc=force/mass,time=5 seconds, v=u+at, d=ut+.5*at^2

OpenStudy (anonymous):

Also note that from the definition of impulse we can define the following expression\[F \cdot t = \Delta p\]where p is the momentum. \[p = m \cdot v\]